OpenStudy (anonymous):

Find the distance between the points given. (0, 6) and (5, 12) √(61) √(11)

OpenStudy (anonymous):

sqrt (11)?

OpenStudy (alexandervonhumboldt2):

(0-5)^2+(6-12)^2=c^2 25+36=c^2 c=sqrt61

OpenStudy (anonymous):

oh makes sense thnx